What is the equilibrium pH of an initially 0.700M for a monoprotic acid with a Ka = 7.687E-4
Answer #1

I have provided the solution. Just go through it. I Let the monoprofic acid be-HX 80 et dissociates as HX -H + + x Tretially 0 0 1 0 out 0 0 at fro at of = teq (0.7-2M RM RM ( XH geven 7 ka = 7.688 1-4 or tota 768X1042 07-27 ² (7.68 x 10) (0.7) 2 (7.687 104 ) * ** + 4.68x10x = ) 53.46X10% =0 2 -768x1P = 3.64x10^M = [H] a pH = -log CH] = -109 (3.64x104) pH = 3.43 AM